OpenStudy (anonymous):

Need major help with Trig. Find the variable Tan(3Beta-4Degrees) = 1/cot(5Beta-8Degrees)

OpenStudy (loser66):

\[tan(3\beta -4)=\frac{1}{cot(5\beta -8)}\]right?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

do you know tan = 1/cot?

OpenStudy (anonymous):

Oh reciprocal identity?

OpenStudy (loser66):

yes, but I just talk about the right hand side. what do you get if you apply that formula? from the right hand side only.

OpenStudy (anonymous):

Uhmm I don't understand how to do that >.<

OpenStudy (loser66):

\[\frac{1}{cot (5\beta -8)}= tan (5\beta -8)\] right?

OpenStudy (anonymous):

yes

OpenStudy (loser66):